If csc of theta is 3 and cos of theta is less than 0, find sin of theta, cos of theta, tan of theta, cot of theta and sec of theta
Answer by stanbon(75887)

If csc of theta is 3 and cos of theta is less than 0,
----
From that you can say that theta is in Q2 where x is negative and y is positive.
csc = r/y = 3/1 implies that r = 4 and y = 1
---
Solve for "x":
r^2 = x^2 + y^2
4^2 = x^2 + 1
x^2 = 15
x = -sqrt(15) (because x is negative in Q2)
-------
find sin = y/r = 1/3
cos = x/r = -sqrt(15)/4
tan = y/x = -1/sqrt(15)
cot = x/y = -sqrt(15)
sec = r/x = -4/sqrt(15)
=================================
Cheers,
Stan H.